The following is the data of recent refrigerator sales at a local Home Depot store.

Month 1 2 3 4 5
Actual Sales 95 100 80 90 ???

Inputs will be exact numbers.

  1. What is the forecasted sales in month 5 using naive approach.
  2. Please use a 2-month simple moving average method to forecast sales in month 5.
  3. Please use a weighted moving average method, with weights of 0.6 one period ago, 0.3 two periods ago, and 0.1 three periods ago, to forecast sales in month 5.
  4. The store manager uses exponential smoothing with a smoothing constant of 0.5 to forecast future sales. Assume the forecasted sales for month 3 is 90. Using the information given, please help the store manager forecast sales for month 5 with exponential smoothing.
Answer #1

1. Using the 2 month simple moving average method

F(5) = {Actual(3) + Actual(4)}/2

= (80+90)/2

= 85

2. Using the 3 month weighted average method

F(5) = {0.6*Actual(4) + 0.3*Actual(3) + 0.1*Actual(2)}

= 0.6*90 + 0.3*80 + 0.1*100

= 54 + 24 + 10

= 88

3. Using the exponential smoothing method

F(4) = 0.5*Actual(3) + (1-0.5)*F(3)

= 0.5*80 + 0.5*90

= 85

F(5) = 0.5*90 + 0.5*85

= 45 + 42.5

= 87.5

Please provide up votes, if you are satisfied with the response. Comment in case of any issue regarding the same.